OpenStudy (anonymous):

Given 3 is one of the roots of the quadratic equation x^2-7x+k=0.Find the value of k.

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

roots means the value that make the equation =0

OpenStudy (anonymous):

so since you have one root just put it instead of x the solve for k

OpenStudy (anonymous):

(3)^2-7(3)+k=0

OpenStudy (anonymous):

got it. answer will be k=12

OpenStudy (triciaal):

since 3 is a root then when x = 3 y = 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnx @AJ01 @UnkleRhaukus @triciaal

OpenStudy (triciaal):

as shown by @AJ01

OpenStudy (anonymous):

9?

OpenStudy (triciaal):

i think you made a switch the constant is ac the sum is b

OpenStudy (unklerhaukus):

ax^2+bx+c =0 the sum of the roots is -b/a the product of the roots is c/a x^2-7x+k=0 if the roots are 3 and m 3 + m = -(-7)/1 = 7 m = 4 the product of the roots is 3 * 4 = k / 1

OpenStudy (anonymous):

tq @UnkleRhaukus